OpenStudy (anonymous):

simplify (5+x^-1)/(25-x^-2)

OpenStudy (anonymous):

a=5 b=x^-1 a^2-b^2=(a+b)(a-b)

OpenStudy (anonymous):

see if you can use the above, to get started.

OpenStudy (angela210793):

@Fiddle r u saying tht (25-x^-2)=(5-x^-1)*(5+x^-1)?

OpenStudy (anonymous):

answer my question

OpenStudy (anonymous):

yes , do you agree ?

OpenStudy (angela210793):

i thought it might b like tht too but wolfram says smth else http://www.wolframalpha.com/input/?i=%2825-x^-2%29

OpenStudy (anonymous):

(5+1/x)/(25-1/x^2)

OpenStudy (anonymous):

\[x \div(5x-1)\]

OpenStudy (angela210793):

^_^

OpenStudy (anonymous):

So wolfram treats both of the below the same way @angela: (25-x^-2) (5-x^-1)*(5+x^-1)

OpenStudy (angela210793):

ignore my answers today...too hot outside :P

OpenStudy (anonymous):

go to the dark side of the moon ! much cooler over there!

OpenStudy (angela210793):

hahah...will do so :D

OpenStudy (anonymous):

anyway - soccerchic do you need more help ? (I hope we didn't get you more confused)

OpenStudy (anonymous):

Nope, i think i got it thanks for everyone's help :)

OpenStudy (anonymous):

OK, if anything is still unclear - just ask here. If you are done with this question - don't forget to click the "Good Answer"s buttons :)

OpenStudy (anonymous):

\[\frac{\frac{1}{x}+5}{25-\frac{1}{x^2}}=\frac{1}{5 (5 x-1)}+\frac{1}{5}=\frac{1}{25 x-5}+\frac{1}{5}=\frac{x}{5 x-1}\]

OpenStudy (anonymous):

i could have sworn i answered this already

OpenStudy (anonymous):

Yes, you did, so I gave you a medal.

OpenStudy (anonymous):

\[-3\pm2\sqrt{5}\]

OpenStudy (anonymous):

satellite - they only asked to simplify, not solve for x :)